OpenStudy (patrickweewoos):

Which equation models this problem? Susan paid $48.80 for eight dish towels. What is the unit price, p, per towel?

OpenStudy (patrickweewoos):

@perl

OpenStudy (igreen):

Just divide. 48.8 / 8 = ?

OpenStudy (patrickweewoos):

6.1

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (patrickweewoos):

wow dude thank you so much and thanks 4 the medal can u help with more

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (patrickweewoos):

I green that wasnt the answer

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (igreen):

It wasn't??

OpenStudy (igreen):

Oh, you have to write an equation..

OpenStudy (patrickweewoos):

A.)8p=48.80 B.)48.80p=8 C.)p/8=48.80 D.)p/48.80=8

OpenStudy (igreen):

It's A because you have to divide 8 to both sides.. \(8p = 48.80\) \(p = \dfrac{48.8}{8}\) Divide: \(p = 6.1\)

OpenStudy (patrickweewoos):

Ok can we do more